You have $ 95 000 to invest. You choose to put $ 145000 into the market by borrowing $ 50000. a. If the​ risk-free interest rate is 5 % and the market expected return is 12 % what is the expected return of your​ investment? b. If the market volatility is 10 %​, what is the volatility of your​ investment?
Answer #1

Weight % invested in Market = 145,000/95,000 = 1.5263

Weight % invested in Risk Free Asset = -0.5263

So,

Expected Return = 1.5263(0.12) - 0.5263(0.05)

Expected Return = 15.68%

Standard Deviation = (1.5263)(0.10)

Standard Deviation = 15.26%